OpenStudy (anonymous):

Given that sum_{n=1}^{infty} = \frac{ pi^2 }{ 6 } and sum_{n=1}^{infty} = \{ a }{1-a} , \left| a |right|\ < 1 determine an reasonable upper bound to sum_{n=1}^{infty} = \frac{ e^(-n) }{ n } by applying the Cauchy Schwartz inequality. Treat the sequences as infinite dimensional vectors.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} = \frac{ \pi^2 }{ 6 }\] and \[\sum_{n=1}^{\infty} = \frac{ a }{1-a} \] 1 determine an reasonable upper bound to\[\sum_{n=1}^{\infty} \frac{ e^{-n} }{ n } \] we cant read that eduation...is this what u wrote?

OpenStudy (anonymous):

it doesnt make sense

OpenStudy (anonymous):

must be |a|<1 instead of 1

OpenStudy (anonymous):

i dont know how i can write that signs

OpenStudy (anonymous):

It doesn't make sense because \(\displaystyle \sum_{n=1}^\infty\) doesn't make sense alone. The infinite sum of what?